OpenStudy (anonymous):

Which container holds the most A. cylinder with a radius 9 in and a height 11 in. B. Sphere with a radius 9 in. C. cone with a radius 12 in. and a height of 20 in.

OpenStudy (anonymous):

for this question you must calculate the volume A) V=Bh=pi(9^2)(11)=891pi B) V=(4/3)pi(r^3)=pi(4/3)(9^3)=972pi C) V=(1/3)Bh=(1/3)(12^2)pi(20)=960pi Therefore the answer is B